OpenStudy (anonymous):

How many miles are equivalent to 89,760 feet? A. 17 miles B. 170 miles C. 1.7 miles D. 1700 miles

OpenStudy (anonymous):

will medal and fan!!!!!!!!!!!!!

OpenStudy (gorv):

i mile =5280 feet

OpenStudy (gorv):

so 5280feet= 1mile 1 feet =1/5280 mile 89760 feet = 89760/5280

OpenStudy (gorv):

17

OpenStudy (anonymous):

thank you :)